OpenStudy (soahelp):

is the answer C If a sprinkler waters 1 over 12 of a lawn in 1 over 2 hour, how much time will it take to water the entire lawn? 7 over 12 hour 1 over 6 hour 6 hours 10 hours

OpenStudy (phi):

same problem as before. can you write the ratio ?

OpenStudy (soahelp):

its not the same problem as before

OpenStudy (soahelp):

so im right that it will take 6 hours to water the full lawn

OpenStudy (phi):

you could do it this way 1/12 of the lawn in 1/2 hour 1/12 in the next 1/2 hour add up to get 2/12 in 1 hour or 1/6 of the lawn in 1 hour how many hours to get all of the lawn? 1/6+1/6+1/6+1/6+1/6+1/6 (it takes an hour for each 1/6)

OpenStudy (soahelp):

im just asking for a simple question and its am i right or no

OpenStudy (phi):

but the faster way is to say \[ \frac{ \frac{1}{2} \ hour}{ \frac{1}{12} \ lawn } = \frac{x \ hours}{1 \ lawn }\]

OpenStudy (soahelp):

x=6